OpenStudy (anonymous):

Jenny must choose a number between61 and 107 that is a multiple of 4, 5, and 10 . Write all the numbers that she could choose

OpenStudy (anonymous):

Looking at the selection of number 4,5,10 what is the least common multiples of these numbers? 4 = 2*2 and 5 = 5*1 10 = 5*2 So it is safe to assume that 5 and 2 will be the limiting values in finding our number between 61 and 107. We know that 5 only shows up evenly at every 10 increments. i.e even 10,20,30 etc. 5 is odd at every increment of 5 5,15,25,35,45.... We also know that 4 nor 10 divide into a whole number, in odd increments of 5. 30/4 , 50/4 , 70/4 , 90/4 are all not whole numbers. 25/10, 35/10, 45/10, etc. So These numbers will be in the actual range of 70-100, and be an even increment of 10. 80 and 100 both lie in the interval of 61 and 107. and both are divisible by 4, 5, and 10.